# Excitable quantum systems: the bosonic avalanche laser

### v2

@@ Abstract
Coherence resonance from quantum stochastic many-body dynamics.—Given the ability of our lasing system to support self-sustained excitations at the mean-field level, the question remains whether or not such a behavior prevails in the presence of noise and when the full many-body dynamics is taken into account. In the current setting, fluctuations arise intrinsically from the discrete nature of the bosonic current, which generates bosonic shot-noise at injection as well as during transport Minoguchi et al. (2025). In the quantum regime with low photon numbers, these fluctuations may either wash out the pulsing cycles or be converted into a regular output signal via the effect of CR. To address this question, we use a stochastic unraveling of Eq. (1) together with Monte-Carlo techniques for an exact simulation of the population dynamics of the gain and the cavity modes, see Sup . In Fig. 4(a) we display sample trajectories of the cavity population in the three lasing phases predicted by our mean-field analysis. All cases show the dominating effect of quantum fluctuations, and there is no longer a clear separation between the different lasing regimes. However, we still observe the emission of semi-regular photon bursts.
@@
Figure 4: (a) Typical stochastic trajectories of the cavity photon number $n_{c}$ for the same parameters as the mean-field results shown in Fig. 2. (b) Plot of the noise spectrum for different values of $\gamma_{\textsl{g}}$, which characterizes the strength of the intrinsic bosonic shot-noise. (c) The coherence parameter $\beta$ given in Eq. (6) is plotted as a function of $\gamma_{\textsl{g}}/\kappa_{c}$, where the stars represent the numerically evaluated values and the solid line is a guide to the eye. A maximum of $\beta$ is found around $\gamma_{\textsl{g}}/\kappa_{c}\approx 1$. For all plots, we have assumed $N=10$, $\kappa_{\ell}/\Gamma=20$, and $\zeta=1/2$.
To quantify this behavior, we evaluate the correlation function
@@
where $\langle\bullet\rangle$ denotes the trajectory average. The shape of the resulting noise spectra is plotted in Fig. 4(b) for different $\gamma_{\textsl{g}}$, which quantifies the strength of the injected noise. We see that for all parameters, the spectrum reaches a maximum, $S(\omega_{\rm max})$, at a nonvanishing frequency $\omega_{\rm max}>0$, which indicates a semi-regular behavior. Deep within the self-pulsing phase, the peak frequency $\omega_{\rm max}\approx 2\pi/\tau$ is roughly consistent with the mean-field prediction for the period $\tau$, although we do not find a tight correspondence in general.
In the literature on excitable systems, it is common to introduce the so-called coherence parameter Gang et al. (1993); Kato and Nakao (2021)
@@
where $\Delta\omega$ is the half-width at half-maximum of the spectral peak. In essence, the parameter $\beta$ quantifies the regularity of the signal produced and a characteristic feature of excitable systems is the existence of a maximum of the coherence parameter as a function of the applied noise strength. For the current system, $\beta$ is plotted in Fig. 4(c) for varying $\gamma_{\textsl{g}}$, from which we make two important observations. First, the coherence parameter changes smoothly across the phase boundaries shown in Fig. 2, which demonstrates that nonlinear features of classical excitable systems do not necessarily translate into the quantum regime. However, $\beta$ still exhibits the characteristic maximum, which motivates the notion of an excitable quantum system in the present case. Here, we have focused on the case $\zeta=1/2$, ie, driving by an infinite-temperature bath, to show that even such a completely random input can be transformed into a regular output. In the SM, we also present results for other, more regular driving forces (in particular pure gain, for $\zeta=1$) Sup . These more regular inputs then lead to an even more marked coherence resonance effect.
@@
Photon-avalanche detection of microwave photons.—Beyond its fundamental interest, the ability to convert even a few injected gain bosons into an avalanche of photons at the cavity output makes this mechanism very promising for quantum sensing and amplification applications. Here, we outline a potential realization of our model in Eq. (1) with a superconducting circuit depicted in Fig. 5 and illustrate its use for microwave photon detection. In this circuit, the cavity mode (shaded in green) as well as the ladder modes (shaded in blue) are represented by quantized $LC$ resonators with frequencies $\omega_{c}$ and $\omega_{p}$ in the microwave regime. Neighboring modes $a_{p}$ and $a_{p+1}$ are coupled to the cavity mode via an effective interaction of the form
@@
which involves an additional dissipative waste mode (shaded in red) with frequency $\omega_{b}$ and annihilation operator $b_{p}$. Photons created in this mode quickly decay with a rate $\kappa_{b}\gg g$ and can thus be adiabatically eliminated to obtain the incoherent hopping terms in Eq. (1) with a rate $\Gamma\simeq 4g^{2}/\kappa_{b}$. In superconducting circuits, multi-photon interactions as in Eq. (7) appear naturally in nonlinear Josephson junctions, and can be selected among other contributions by enforcing the resonance condition $\omega_{p}-\omega_{p+1}+\omega_{b}+\omega_{c}=\omega_{e}$, where $\omega_{e}$ is the modulation frequency of an externally applied magnetic flux.
@@
Figure 5: (a) Sketch of a superconducting circuit layout for realizing a bosonic avalanche laser. A set of LCLC resonators representing both the common cavity mode (green) and the ladder modes (blue) are coupled via nonlinear SNAIL-type Frattini et al. (2017) couplers (spiral) to an additional set of strongly damped waste modes (red). As discussed in End Matter, the coupling can be modulated by external fluxes $\phi_{e}(t)$ to achieve a resonant four-mode interaction of the form given in Eq. (7). After eliminating the waste modes, we obtain the dissipative three-mode hopping in Eq. (1). (b) Sample trajectories showing the cavity population $n_{c}$ over time for different initial populations $n_{1}$ of the first ladder mode. For the same conditions, (c) shows the corresponding distribution of the integrated output signal of the cavity mode, $n_{\rm out}$, for $500$ trajectories. For these plots, we have assumed $N=10$, $\kappa_{\ell}=10\Gamma$, $\kappa_{c}=0.2\Gamma$ and also that each ladder mode decays with a rate $\kappa_{0}=0.2\Gamma$.
In End Matter we present a more detailed analysis of this circuit and the conditions that are required to obtain Eq. (1). This analysis shows that for state-of-the-art experimental parameters, the realization of a microwave avalanche laser with N=510N=5-10 ladder modes and a hopping rate of $\Gamma/(2\pi)\approx 100$ kHz can be achieved. This rate exceeds the intrinsic loss rate $\kappa_{0}/(2\pi)=10-100$ kHz of high-Q microwave resonators Frunzio et al. (2005); Reagor et al. (2013); Somoroff et al. (2023)f, such that individual photons can still hop through the ladder of ladder modes, before they decay. For these conditions, we simulate the population of the cavity mode under the assumption that the first ladder mode is initialized with a well-defined number of photons, $n_{1}=1,\dots,5$. In Fig. 5(b) and (c) we plot individual trajectories of $n_{c}(t)$ as well as a histogram of the integrated output signal, $n_{\rm out}=\int_{0}^{T}dt\,I_{c}(t)$, where $I_{c}(t)$ is the photon current emitted by the cavity. We see that the initial photon-number states are both amplified and separated according to the initial state. Therefore, when combined with a regular microwave amplifier, capable of resolving those amplified wavepackets, this device realizes a number-resolved detector for microwave photons.
@@
Conclusion.—In summary, we have analyzed the behavior of a bosonic avalanche laser, where a dissipative three-mode mixing process converts a bosonic input current into an amplified signal at the cavity output. We have shown that in this system, mutually stimulated processes can lead to a self-pulsing phase at the mean-field level, which survives as a CR even deep in the quantum regime, where the dynamics is dominated by bosonic shot-noise. This conversion of quantum fluctuations into an amplified, semi-regular signal can potentially be used in quantum sensing applications, and we described the implementation of a number-resolved microwave photon detector as a specific example. The underlying mechanism and the described behavior can further be of relevance for the development of autonomous quantum engines and clocks Mari et al. (2015); Serra-Garcia et al. (2016); Carollo et al. (2020); Antonio Marín Guzmán et al. (2024), where the interplay between periodic motion and quantum fluctuations is at the heart of the subject and still not well understood.

### v3

@@ III Quantum stochastic dynamics
Given the ability of our lasing system to support self-sustained excitations at the mean-field level, the question remains whether or not such a behavior prevails in the presence of noise and when the full many-body dynamics is taken into account. In the current setting, fluctuations arise intrinsically from the discrete nature of the bosonic current, which generates bosonic shot-noise at injection and during transport [25]. In the quantum regime with low photon numbers, these fluctuations are expected to wash out the pulsing cycles. In excitable systems, however, the mechanisms of SR and CR can also induce the opposite effect and convert the noisy input into regular output signals.
To investigate the role of these competing effects on the overall behavior of our lasing system, we use a stochastic unraveling of Eq. (1) together with Monte-Carlo techniques to perform exact simulations of the population dynamics of the gain and the cavity modes (see Appendix D for the details of the numerical method). Note that the noise process used in those simulations depends on the pumping mechanism defined by $\mathcal{L}_{\rm gain}$ in Eq. (3). For now, we focus on the infinite-temperature limit $\zeta=1/2$, for which—for a given average pump rate—fluctuations are maximal. In our simulation, this process is implemented by randomly injecting particles in the first chain site at a rate $\gamma_{\rm inj}=\gamma_{\textsl{g}}(1+n_{1})$, and extracting them at a rate $\gamma_{\rm extr}=\gamma_{\textsl{g}}n_{1}$, with $n_{1}$ the number of particles already present on the first site. The average gain rate is then given by $\gamma_{\rm inj}-\gamma_{\rm ext}=\gamma_{\textsl{g}}$. We return to a discussion about other types of injection processes in Sec. III.3 below.
In Fig. 4(a), we display sample trajectories of the cavity population in the three phases predicted by our mean-field analysis. All cases show the dominating effect of quantum fluctuations, and there is no longer a clear separation between the different lasing regimes. However, we still observe the emission of semi-regular photon bursts. To quantify this behavior, we follow two different approaches and analyze, on the one hand, the noise spectrum of the output field, and on the other hand, the inter-spikes interval statistics.
@@
Figure 4: (a) Typical stochastic trajectories of the cavity photon number $n_{c}$ (same parameters as the mean-field results shown in Fig. 2). (b) Plot of the noise spectrum for different values of $\gamma_{\textsl{g}}$, which characterizes both the average pumping rate and the strength of the intrinsic bosonic shot-noise. (c) The coherence parameter $\beta$ defined in Eq. (8) is plotted as a function of $\gamma_{\textsl{g}}/\kappa_{c}$, where the stars represent the numerically evaluated values and the solid line is a guide to the eye. A maximum of $\beta$ is found around $\gamma_{\textsl{g}}/\kappa_{c}\approx 1$. For all plots, we have assumed $N=10$, $\kappa_{\ell}/\Gamma=20$, and $\zeta=1/2$.
III.1 Noise spectrum and coherence parameter
First, we evaluate the correlation function
@@ III.1 Noise spectrum and coherence parameter
where $\langle\bullet\rangle$ denotes the trajectory average. The shape of the resulting noise spectra is plotted in Fig. 4(b) for different $\gamma_{\textsl{g}}$, which quantifies the strength of the driving process and of the injected noise. We see that for all parameters, the spectrum reaches a maximum, $S(\omega_{\rm max})$, at a nonvanishing frequency $\omega_{\rm max}>0$, which indicates a semi-regular behavior. Deep within the self-pulsing phase, the peak frequency $\omega_{\rm max}\approx 2\pi/\tau$ is roughly consistent with the mean-field prediction for the period $\tau$, although we do not find a tight correspondence in general.
In the literature on excitable systems, it is common to introduce the so-called coherence parameter [6, 17]
@@
where $\Delta\omega$ is the half-width at half-maximum of the spectral peak. In essence, the parameter $\beta$ quantifies the regularity of the signal produced; a characteristic feature of excitable systems is the existence of a maximum of the coherence parameter as a function of the applied noise strength. For the current system, $\beta$ is plotted in Fig. 4(c) for varying $\gamma_{\textsl{g}}$, from which we make two important observations. First, the coherence parameter changes smoothly across the phase boundaries shown in Fig. 2, which demonstrates that nonlinear features of classical excitable systems do not necessarily translate into the quantum regime. However, $\beta$ still exhibits the characteristic maximum, which is reminiscent of the behavior found in excitable systems exhibiting CR. We will discuss this connection further below.
The results presented in Fig. 4, which are obtained for the limit of an infinite-temperature bosonic reservoir with $\zeta=1/2$, show that even this maximally noisy input is converted into a semi-regular output. In Appendix E, we also present simulation results for a pure gain process with $\zeta=1$, for which we observe an even more pronounced maximum of the coherence parameter.
III.2 Inter-spike interval statistics
To consolidate these findings, we also conducted a direct analysis of the inter-burst intervals. Contrary to other excitable systems [2], the pulses in our setup present fairly large fluctuations in both height and emission time. Thus, to identify a burst, we only take into account signals that exceed a certain threshold (here set at twice the time-average of the signal). We then fit the part of each trajectory that lies above this threshold by Lorentzian functions, in order to remove artifacts coming from short-time oscillations and multiple subpeaks. This fitting procedure is illustrated for a few sample trajectories in Fig. 5(a). From the centers of the Lorentzians, we obtain the set of burst times $t_{\rm b}^{(i)}$, from which we sample the distribution of interval spacings $\tau_{i}=t_{\rm b}^{(i)}-t_{\rm b}^{(i-1)}$.
The resulting distributions for the interval spacings are presented in Fig. 5(b) for different values of $\gamma_{\rm g}$. As we increase the pumping strength, the interval distribution shifts towards smaller values, while its variance tends to decrease. This is in agreement with the analysis in frequency space, which showed noise spectra acquiring a larger mean and variance for increasing pumping strength. Furthermore, the maximum of the distribution, although it deviates from the mean-field prediction, largely coincides with the corresponding maximum in the noise spectrum. To quantify the regularity of the distribution, we plot in Fig. 5(c) the ratio between the mean interval spacing and its standard deviation, $\langle\tau\rangle/\sqrt{\langle\tau^{2}\rangle-\langle\tau\rangle^{2}}$. This quantity exhibits a maximum around $\gamma_{\textsl{g}}/\kappa_{c}=1$, again in line with the behavior in the frequency domain.
Figure 5: Inter-spike interval statistics. (a) Illustration of the fitting procedure used to identify pulses in the output of the lasing cavity. The actual trajectories obtained from Monte-Carlo simulations (blue lines) are overlapped with the fitted Lorentzian curves (red lines), whose maxima lie at times $t_{\rm b}^{(i)}$. For these examples, we have assumed the ratios $\gamma_{\rm g}/\kappa_{c}=0.3,0.6,1$ (from top to bottom). (b) Histogram of the intervals $\tau_{i}$ between successive bursts, for the same three $\gamma_{\rm g}/\kappa_{c}$ values.(c) Plot of the ratio between the mean burst interval and its standard deviation as a function of the pump strength. All other parameters in the plots are the same as in Fig. 4.
III.3 Role of the noise
In the numerical simulations presented above, we have observed a maximum for the coherence parameters in both frequency and time space. Both quantities indicate a certain amount of regularity of the emitted pulses, in spite of an overall strongly fluctuating behavior. These maxima are reached as a function of the gain rate $\gamma_{\rm g}$, which, in view of the shot noise of the injected particle current, is also a measure for the amount of noise that is injected into the system. The observed behavior is thus reminiscent of CR in classical excitable systems, where similar maxima in the coherence parameter occur. CR, however, is usually associated with a quasi-regular response of a noise-driven system, which shows no particular dynamics in the absence of the noise. Here, instead, self-oscillations may already occur at the mean-field level.
Leaving aside the precise correspondence between the observed quasi-regular pulses and SR and CR in the conventional sense, it would be interesting to see more directly the influence of noise on the production of a semi-regular output. In the current model, this influence is somewhat hidden by the fact that the ladder is driven by an average, i.e., deterministic current, with shot-noise fluctuations on top. To separate those two contributions, we performed additional numerical simulations, where we replaced the gain process in Eq. (3) by
$\mathcal{L}_{\rm gain}\,\,\rightarrow\,\,x\mathcal{L}_{\rm gain}+y\mathcal{L}_{\rm Poisson}+(1-x-y)\mathcal{L}_{\rm reg},$ (9)
where $0\leq x+y\leq 1$. Here, $\mathcal{L}_{\rm reg}$ and $\mathcal{L}_{\rm Poisson}$ both describe incoherent processes in which bosons are injected one by one into mode $p=1$, independently of the prior occupation of the mode. This is to be contrasted with $\mathcal{L}_{\rm gain}$, in which driving is achieved by a balance of gain and loss (see discussion at the beginning of Sec. III). For $\mathcal{L}_{\rm reg}$, the injection occurs at quasi-regular times, spaced by $\gamma_{\textsl{g}}^{-1}$  [^1] More precisely, the probability that an injection takes place in the interval $\tau,\tau+d\tau]$ obeys a Gaussian distribution: $p_{e}(\tau,\tau+d\tau)=\frac{2}{\Delta\sqrt{\pi}}e^{-\left(\frac{\tau-\tau_{0}}{\Delta}\right)^{2}}dt$. In the limit $\Delta\rightarrow 0$, particles are injected exactly at times $\tau_{n}=n\tau_{0}$, thereby realizing a deterministic injection at a rate $1/\tau_{0}=\gamma_{\textsl{g}}$. In practice, we found that values of $\Delta\leq\gamma_{\textsl{g}}$ were enough to achieve convergence.. For $\mathcal{L}_{\rm gain}$, by contrast, these intervals follow a random Poisson distribution, again with average $\gamma_{\textsl{g}}^{-1}$ [^2] At the Lindbladian level, this would be realized by a jump operator: $\mathcal{L}_{\text{gain}}=\mathcal{D}\left[\frac{1}{\sqrt{1+a^{+}a}}a^{+}\right]$. This means that, for $x=1,y=0$ or $x=0,y=1$, we recover a noisy drive, respectively through an infinite-temperature reservoir or a particle-by-particle Poisson process; whereas, for $x=y=0$, the injection becomes quasi-deterministic. By varying $x$ and $y$, this model allows us to tune the level of randomness in the gain process, while keeping the same average injection rate. Note that the process $\mathcal{L}_{\rm Poisson}$ may actually be quite difficult to engineer experimentally, but its theoretical appeal is to offer a direct counterpart to the deterministic process; since these two processes differ only by the injection interval distribution, any difference between them can be fully attributed to the stochastic nature of the drive.
In Fig. 6, we show that a larger $x$ or $y$ (therefore, a noisier process) can significantly improve the coherence of the output. For the parameters displayed here, there is even an optimal value of $x$ for which the coherence is maximized, which is again suggestive of CR. By isolating the contribution of random fluctuations in the drive, this analysis thus proves that noise does help in creating a more coherent output, showing how this key feature of excited systems is carried out in the quantum regime.
In general, the problem of separating the stochastic and deterministic contributions in the drive is not straightforward when dealing with weakly-driven systems in the quantum regime: we discuss this issue in more details in Appendix F. Beyond the specific solution we have adopted to characterize the system at hand, a more systematic approach may be needed to study further examples of quantum excitable systems. We leave this point to future investigations.
Figure 6: (a) Pumping scheme of Eq. (9), for $x=0$. Particles are injected one by one at random intervals $\tau$, whose distribution $p(\tau)$ is the sum of a stochastic and a quasi-deterministic process. The limits $y\rightarrow 1$ and $y\rightarrow 0$ give respectively fully random and fully periodic injection events. For $x\neq 0$, we have instead a balance between injection and extraction processes. (b) Coherence parameter versus stochasticity of the drive, for a constant total rate $\gamma_{\textsl{g}}=0.6\kappa_{c}$. The quasi-deterministic process $\mathcal{L}_{\rm reg}$ is compared to an infinite-temperature bath (top) or a pure Poisson process (bottom). In both cases, the added stochasticity contributes positively to the regularity of the output.
IV Applications: An avalanche detector for microwave photons
@@
Beyond its fundamental interest, the ability to convert even a few injected gain bosons into an avalanche of photons at the cavity output makes this mechanism very promising for quantum sensing and amplification applications. As a specific use-case, we discuss in this section the realization of our model in Eq. (1) with superconducting circuits, and illustrate its use for microwave photon detection.
IV.1 Circuit QED implementation: general approach
To realize the cascaded lasing process in the microwave regime, we consider a superconducting circuit as outlined in Fig. 7(a). In this circuit, the extended cavity mode (shaded in green) as well as the individual ladder modes (shaded in blue) are represented by quantized $LC$ resonators with frequencies $\omega_{c}$ and $\omega_{p}$ in the microwave regime. To induce the targeted nonlinear dissipation processes, each pair of neighboring ladder modes is coupled to the cavity and an additional dissipative waste mode (shaded in orange), with frequency $\omega_{b}$ and annihilation operator $b_{p}$, through a nonlinear Josephson coupler. This coupling element generates interaction terms at all orders between the various modes. Here we are interested in fourth-order contributions of the form
@@ IV.1 Circuit QED implementation: general approach
which can be selected among other fourth-order and lower-order terms by driving the system with an externally applied magnetic flux at a frequency $\omega_{e}=\omega_{p}-\omega_{p+1}+\omega_{b}+\omega_{c}$. This coupling generates the process $a_{p}a^{\dagger}_{p+1}c^{\dagger}$ that we are interested in, plus the creation of an additional excitation in the waste mode. Photons created in this mode quickly decay with a rate $\kappa_{b}\gg g$, and can thus be adiabatically eliminated to obtain the incoherent hopping terms in Eq. (1), with a rate $\Gamma\simeq 4g^{2}/\kappa_{b}$. The full model is then achieved by repeating this unit cell in a chain configuration.
Following this general approach, we present in Appendix G a more detailed discussion of the individual circuit elements and the main assumptions that are necessary to arrive at the process given in Eq. (10) and the resulting master equation in Eq. (1). This analysis shows that, for state-of-the-art experimental parameters, the realization of a bosonic avalanche laser with N=510N=5-10 ladder modes and a hopping rate of $\Gamma/(2\pi)\approx 100$ kHz can be achieved. This rate exceeds the intrinsic loss rate $\kappa_{0}/(2\pi)=10-100$ kHz of high-Q microwave resonators [5, 28, 31], such that individual photons can still hop through the whole ladder before they decay.
@@
Figure 7: (a) Sketch of a superconducting circuit layout for realizing a bosonic avalanche laser. A set of LCLC resonators representing both the common cavity mode (green) and the ladder modes (blue) are coupled via nonlinear SNAIL-type [4] couplers (spiral) to an additional set of strongly damped waste modes (orange). As discussed in Appendix G, the coupling can be modulated by external fluxes $\phi_{e}(t)$ to achieve a resonant four-mode interaction of the form given in Eq. (10). After eliminating the waste modes, we obtain the dissipative three-mode hopping in Eq. (1). (b) Sample trajectories showing the cavity population $n_{c}$ over time for different initial populations $n_{1}$ of the first ladder mode. For the same conditions, (c) shows the corresponding distribution of the integrated output signal of the cavity mode, $n_{\rm out}$, for $500$ trajectories. For these plots, we have assumed $N=10$, $\kappa_{\ell}=10\Gamma$, $\kappa_{c}=0.2\Gamma$, and also that each ladder mode decays with a rate $\kappa_{0}=0.2\Gamma$.
IV.2 Microwave Photon Detection
Based on the parameters estimated for the circuit QED implementation above, we simulate the cavity population under the assumption that the first ladder mode is initialized with a well-defined number of photons, $n_{1}=1,\dots,5$. In Fig. 7(b), we first show the successive evolution of the cavity population for individual trajectories of $n_{c}(t)$. We see that already a single photon can trigger a large excitation of the lasing cavity mode, despite significant losses in each of the ladder modes. As the initial number of photons increases, the build-up of this excitation is both accelerated and amplified, consistent with the mechanism described in Fig. 3 for large photon numbers. For the same parameters, Fig. 7(c) shows a histogram of the integrated output signal,
$n_{\rm out}=\int_{0}^{T}dt\,I_{c}(t),$ (11)
where $I_{c}(t)=\kappa_{c}n_{c}(t)$ is the photon current emitted by the cavity. We see that different initial photon-number states give rise to output signals that are well separated, approximately by a photon number corresponding to the number of ladder modes. The residual overlap between neighboring distributions can be attributed to the loss of a photon in the initial mode, $p=1$, before the accelerated cascade process has started. Losses at a later stage are less relevant. Therefore, when combined with a regular microwave amplifier capable of resolving those amplified wavepackets, this device realizes a number-resolved detector for microwave photons.
V Conclusion
@@
In summary, we have analyzed the behavior of a bosonic avalanche laser, in which a dissipative three-mode mixing process converts a bosonic input current into an amplified signal at the cavity output. We have shown that in this system, mutually stimulated processes can lead to a self-pulsing phase at the mean-field level, which survives as a train of noisy but semi-regular output bursts even deep in the quantum regime, where the dynamics is dominated by bosonic shot-noise. Interestingly, the regularity of these bursts can be enhanced by increasing the amount of noise, and we observed a maximum of the coherence parameter, reminiscent of CR in conventional excitable systems. This conversion of quantum fluctuations into an amplified, semi-regular signal can potentially be used in quantum sensing applications, and we described the implementation of a number-resolved microwave photon detector as a specific example. The underlying mechanism and the described behavior can also be of relevance for the development of autonomous quantum engines and clocks [24, 30, 3, 1], where the interplay between periodic motion and quantum fluctuations is at the heart of the subject and still not well understood.
Acknowledgements.